### v1

@@
Tidal structures around globular clusters provide valuable insights into cluster evolution and the hierarchical assembly of the Milky Way. Using wide-field imaging data from the DESI Legacy Survey combined with a color-magnitude matched-filter technique, we perform a systematic analysis of extra-tidal features in 28 Galactic globular clusters of likely extragalactic origin, representing the largest homogeneous sample examined in this context to date. The clusters display diverse morphologies: 12 exhibit tidal tails, 9 show diffuse envelopes, and 7 reveal no clear extra-tidal features. Notably, we report the first detection of an extended tidal structure around the Sagittarius-associated cluster Terzan 7. To explore the underlying drivers, we compare intrinsic properties, orbital dynamics, and possible accretion associations across morphological groups. From the parameter distributions, complemented by Kolmogorov-Smirnov tests, we find that total mass, escape velocity, concentration, tidal filling factor, pericentric radius, eccentricity and radial angle in action-angle coordinates are all likely group-sensitive parameters. These results suggest that both internal cluster properties and orbital configurations play important roles in shaping extra-tidal morphologies. In addition, the cluster’s accretion history shows no clear correlation with the presence of tidal features, indicating that it is not a direct driver of outer structure formation. Overall, the diversity of tidal structures is unlikely to be governed by a single factor, but instead reflects the interplay between internal dynamical evolution and the external Galactic environment. This study provides the most comprehensive constraints so far on the physical processes driving extra-tidal structures in accreted globular clusters.
@@ 1 Introduction
Tidal structures are generally believed to form when stars acquire sufficient energy to escape the gravitational potential of their host globular clusters. In the early stages of cluster evolution, energetic feedback from stellar evolution, gas expulsion, and dynamical disequilibrium can cause young clusters to rapidly lose their gravitational binding and dissolve (2001MNRAS.323..988G; 2001MNRAS.322..231K; 2006MNRAS.373..752G; 2007MNRAS.380.1589B). Over longer timescales, stellar escape can be driven by a combination of internal and external mechanisms. Internally, two-body relaxation redistributes energy among stars, enabling some to reach escape velocity (2008MNRAS.389..889K; 2008MNRAS.387.1248K). Externally, tidal stripping induced by the Galactic potential, as well as dynamical perturbations such as disk and bulge shocking during the cluster’s orbital passages through the Galactic plane or central regions, can further enhance stellar mass loss and facilitate the development of extratidal features(1997MNRAS.289..898V; 2003MNRAS.340..227B; 2007ApJ...659.1212M; 2015MNRAS.446.3100H).
Due to the complex formation pathways and the diverse Galactic environments in which they reside, globular clusters exhibit a wide variety of tidal structures, such as symmetric or multi-armed tidal tails, and diffuse outer envelopes. However, the physical mechanisms driving the formation of these different tidal structures are still under active investigation. Early numerical N-body simulations by 2007ApJ...659.1212M suggest that repeated pericentric passages or strong interactions with the dense regions of the Galaxy can trigger the stripping of outer stars, leading to the formation of tidal tails. In contrast, the work by 2010MNRAS.401..105K; 2012MNRAS.420.2700K highlights the role of epicyclic motions of escaping stars in shaping the formation and morphology of tidal features. Other studies have pointed out that orbital parameters, such as inclination and eccentricity, play an important role in regulating the mass loss of clusters (2019ApJ...882...98P; 2020A&A...637L...2P), which in turn affects their susceptibility to developing tidal structures. Meanwhile, the potential connection between cluster origin and structural properties has garnered increasing attention. For example, whether the presence of tidal structures is connected to specific accretion events, such as the Sagittarius dwarf galaxy or the Gaia-Sausage-Enceladus merger, remains to be systematically investigated.
Benefiting from the advent of wide-field imaging surveys such as the Sloan Digital Sky Survey (SDSS; 2009ApJS..182..543A), Pan-STARRS (2016arXiv161205560C), the Dark Energy Survey (DES; 2018ApJS..239...18A), and the DESI Legacy Imaging Surveys (2019AJ....157..168D), together with the development of various techniques optimized for identifying tidal substructures, most notably the application of matched filter method, previous studies have uncovered an increasingly rich diversity of tidal features, including long tidal tails (e.g., Palomar 5 and NGC 5466; 2006ApJ...641L..37G; 2006ApJ...639L..17G), multi-arm structures (e.g., NGC 4147; 2010A&A...522A..71J; 2024AJ....168..237Z), as well as a variety of other complex substructures reported in the literature (2006ApJ...651L..33L; 2009AJ....138.1570O; 2010AJ....139..606C; 2010A&A...522A..71J; 2012MNRAS.419...14C; 2014MNRAS.445.2971C; 2016MNRAS.461.3639K; 2017MNRAS.467L..91C; 2017ApJ...841L..23N; 2018MNRAS.474..683C; 2018MNRAS.473.2881K; 2018ApJ...860...66M; 2018MNRAS.473.3062M; 2018ApJ...862..114S; 2019MNRAS.486.1667C; 2019MNRAS.485.4906D; 2019ApJ...884..174G; 2019MNRAS.488.1535P; 2020A&A...635A..93P; 2020AJ....160..244S; 2020MNRAS.495.2222S; 2020ApJ...902...89T; 2021ApJ...914..123I; 2021A&A...646A.176P; 2022ApJ...930...23N; 2022MNRAS.509.3709P; 2022MNRAS.513.3136Z; 2025A&A...693A..69A; 2025AJ....170..294C; 2025MNRAS.540.2863W; 2025ApJ...988...39W). Precisely because such a large sample of globular clusters with diverse tidal morphologies has now become available, we are afforded the opportunity to investigate, from an observational perspective, the key physical mechanisms that drive the diversity of tidal structures in these systems.
Recent studies have utilized observational data to explore the key factors that drive the formation of different types of tidal structures. 2020A&A...637L...2P conducted a multi-parameter analysis of 53 Galactic globular clusters compiled from previous literature, including both in-situ and accreted systems. However, their classification of tidal structures relied on results from multiple studies that used heterogeneous datasets obtained with different telescopes. More recently, 2025AJ....170..157K constructed a smaller sample of 30 globular clusters from the Pristine-Gaia-Synthetic (PGS) catalogue, likewise covering both in-situ and accreted clusters, and in a brief section of their work also performed an exploratory parameter correlation analysis. These studies provide valuable attempts to explore possible connections between cluster properties and tidal structures, although no significant correlations were found. Nevertheless, studies based on homogeneous data that focus exclusively on a specific class of globular cluster systems, such as those accreted from outside, are still lacking. Such clusters, owing to their external origin, may have undergone distinct pathways of dynamical evolution and tidal disruption (2019A&A...630L...4M; 2020MNRAS.493..847F).
Therefore, the objective of this paper is to investigate the key factors that give rise to the tidal structures of the accreted globular cluster population by employing a homogeneous dataset and a unified methodology. We use the deepest dataset among the currently available public data, and the largest accreted cluster sample among those widely recognized (2019A&A...630L...4M; 2020MNRAS.493..847F). Combined with a consistent matched-filter technique, this study will minimize methodological biases and systematic uncertainties.
This paper is organized as follows. Section 2 presents the selection criteria for the globular cluster sample and introduces the photometric data employed. Section 3 outlines the matched-filter technique adopted to search for tidal tails around globular clusters. Section 4 presents the main results, including the identification and classification of tidal structures, their correlation with physical parameters, orbital properties, and the Galactic environment, as well as a discussion of the limitations in detecting such structures. Section 5 summarizes the main findings of this study.
2 Observational material
We began by constructing an initial list of candidate globular clusters in the Galaxy with probable extragalactic origins, based on consistent classifications in both 2019A&A...630L...4M and 2020MNRAS.493..847F. In the original 2019 version, 2019A&A...630L...4M classified clusters primarily according to their positions in integrals-of-motion (IOM) space derived from Gaia DR2 astrometry, with the age-metallicity relation serving as a secondary criterion. In the 2025 update, this classification was refined using Gaia eDR3 astrometry, together with two independent sets of distance estimates from 2021MNRAS.505.5957B and 1996AJ....112.1487H (1996AJ....112.1487H, 2010 edition). We retained only those clusters that were consistently classified as accreted under both distance assumptions. In contrast, 2020MNRAS.493..847F adopted the IOM framework of 2019A&A...630L...4M but also incorporated the updated compilation of cluster ages and metallicities from 2019MNRAS.486.3180K, increasing the number of clusters with complete chemical information. Where available, they included $\alpha$-element abundances from 2020MNRAS.493.3363H, enabling more reliable assignments of clusters to disrupted dwarf galaxies and, in some cases, revising the original classifications. By cross-matching the results of these two studies and excluding clusters with unclear or ambiguous origins, we obtained an initial sample of 44 globular clusters of extragalactic origin.
[…]
@@ 2 Observational material
We selected stellar sources classified as point-source in the DESI Legacy Surveys catalogs, requiring $g\leq 24$ and $r\leq 23$, with photometric uncertainties less than 0.2 mag in both bands, in order to ensure a more complete and reliable star sample. For clusters lacking $r$-band data (e.g., Ter 7, Ter 8, and Pal 12), we substituted $i$-band photometry and applied the same selection criteria to maintain consistency. In the DESI Legacy Surveys catalogs, some photometric measurements (such as those in the $i$-band) are drawn from supplementary datasets incorporated into the survey products (e.g., the Dark Energy Survey, the DELVE Survey, and the DeROSITA Survey). Subsequently, all magnitudes were corrected for Galactic extinction using the dust maps of 1998ApJ...500..525S, recalibrated by 2011ApJ...737..103S, with extinction coefficients of 3.214, 2.165, 1.592, and 1.211 for the $g$, $r$, $i$, and $z$ bands, respectively, following the calibrations of 2019AJ....157..168D.
@@
Finally, we examined whether each cluster in the initial sample lies within the footprint of the DESI Legacy Imaging Surveys and possesses sufficient photometric coverage. Among them, 11 clusters (NGC 2298, NGC 3201, NGC 4833, NGC 5286, NGC 6235, NGC 6333, IC 1257, FSR 1758, ESO-SC06, NGC 6715, and NGC 6779) are located in regions not covered by DESI Legacy Surveys. In addition, NGC 2808 and NGC 6101 are covered only in the $g$ and $i$ bands, respectively; NGC 5139 suffers from severely incomplete coverage, with many areas lacking photometric detections; IC 4499 is heavily contaminated by field stars and has poor photometric quality; and NGC 2419, at a distance of about 88 kpc (2021MNRAS.505.5957B), is too remote for our photometric selection criteria to be reliably applied. After excluding these systems, we obtained a final working sample of 28 clusters (see Table 1). These include clusters with well-established progenitors (such as Gaia-Sausage-Enceladus (GSE), Sagittarius (Sag), Sequoia (Seq), and the Helmi streams (H99)) as well as clusters of clear accreted origin but uncertain association with a specific progenitor. Although some previous compilation studies have reported a comparable number of accreted globular clusters, the sample constructed in this work is entirely based on uniformly deep imaging data from the DESI Legacy Surveys, making it the deepest and largest homogeneous dataset to date for studying the tidal structures of globular clusters of extragalactic origin within a consistent data framework.
@@
Table 1: Internal parameters and classification of our globular cluster sample. [^1] Columns list: cluster name, King and dynamical tidal radii, present-day cluster mass (logarithmic scale), central escape velocity, central concentration, ratio of half-mass radius to dynamical tidal radius, mass-loss fraction, metallicity, relaxation-age ratio, classification type, and possible progenitor system. Among them, $r_{\rm t}^{\rm King}$, $c$ and [Fe/H] are taken from the Harris catalog (1996AJ....112.1487H 1996AJ....112.1487H, 2010 edition); the classification type and progenitor system are derived in this work; all other parameters are taken from, or computed using, the globular cluster database used in this study.
@@ 3 Matched filter method
To detect potential extra-tidal structures around each globular cluster, we employed the matched filtering technique developed by 2002AJ....124..349R, which enhances the contrast between cluster members and field stars in CMD space and is effective for detecting low-surface-brightness structures. This technique has been successfully applied in several subsequent studies of tidal features around star clusters (e.g., 2010AJ....139..606C; 2010A&A...522A..71J; 2017ApJ...841L..23N; 2018MNRAS.474..683C; 2018ApJ...862..114S; 2020MNRAS.495.2222S; 2022ApJ...930...23N; 2024AJ....168..237Z), demonstrating its effectiveness in tracing surrounding stellar substructures.
According to 2002AJ....124..349R, at any position on the sky (within a solid angle $d\Omega$), the expected number of stars in any sky bin $(i,j)$ in color-magnitude can be expressed as:
@@
Here $f_{\mathrm{cl}}$ represents the normalized color-magnitude distribution of cluster members and serves as the cluster template in the matched-filter method, while $n_{\mathrm{bg}}$ denotes the background number density distribution in the same bin, which is used as the corresponding background template. The parameter $\alpha$ serves as a scaling factor that quantifies the number (or surface density) of stars following the cluster’s CMD distribution within the given sky region. Subsequently, the calculation of this parameter can be derived using the minimum-variance estimator (see Equation 2 in 2002AJ....124..349R).
@@
Based on this model, we first constructed the cluster CMD template ($f_{\mathrm{cl}}$) used in the matched-filtering analysis. To do so, we generated a Hess diagram for the cluster region using bin sizes of $\Delta(g-r)=0.05$ and $\Delta r=0.1$ mag. In most cases, the CMD was built in the (gr,r)(g-r,r) plane. However, for a few clusters lacking $r$-band photometry, such as Ter 7, Ter 8 and Pal 12, the (gi,i)(g-i,i) combination was adopted instead. In order to reduce the effects of crowding and incompleteness in the cluster core, stars within the inner 1′-3′ were excluded from the template construction. For clusters with relatively small tidal radius and more diffuse stellar distributions (e.g., Whiting 1 and Pal 13), no such exclusion was applied. Consequently, the cluster template was constructed from stars located between this exclusion zone (when applied) and the cluster’s tidal radius (from 1996AJ....112.1487H (1996AJ....112.1487H), 2010 edition). In addition, to reduce field-star contamination and obtain a clean cluster CMD template $f_{\mathrm{cl}}$, we selected stars with 0.5<gr<1.0-0.5<g-r<1.0 and $14<r<23$ mag, and subtracted a local background derived from a comparing field in an annular region centered on the cluster, typically located between about $1^{\circ}$ and $1.5^{\circ}$ from the cluster center. The stellar density in this annular region was normalized to the area of the cluster region and subtracted from the cluster Hess diagram, yielding a background-subtracted CMD template (see the upper panels of each Hess diagram in Fig. 1). In contrast, the global background CMD template ($n_{\mathrm{bg}}$) required for the matched-filtering analysis was constructed from four rectangular control regions ($1^{\circ}\times 1^{\circ}$) located at different position angles far away from the cluster, each more than $2^{\circ}$ from the cluster center. The Hess diagrams of these regions were then averaged to produce a uniform global background model (see the lower panels of each Hess diagram in Fig. 1).
@@
Finally, the matched-filter procedure yields the quantity $\alpha$, which represents the number of stars that pass the matched filter within a given spatial bin. Using these $\alpha$ values, we then constructed a spatial density map to reveal possible tidal structures. However, if the spatial bins are too large, thin tidal features may be smeared out, whereas overly small bins introduce noise and may generate spurious features. Therefore, after testing a range of spatial binning scales and visually inspecting the results, we ultimately adopted a uniform grid with an angular resolution of $0.02^{\circ}\times 0.02^{\circ}$. To evaluate the statistical significance of the detected structures, we computed a signal-to-noise (S/N) map. For each spatial bin, we subtracted the mean background level (estimated from the outer regions excluding the cluster) from its $\alpha$ value and then divided the result by the standard deviation of the background. This final S/N map reveals potential extra-tidal features with enhanced contrast (see Fig. 3).
@@
Figure 1: CMDs for our cluster sample. Each panel displays the template Hess diagram (top), and the global background Hess diagram (bottom).
Figure 2: Continued.

### v2

@@
Tidal structures around globular clusters provide valuable insights into cluster evolution and the hierarchical assembly of the Milky Way. Using wide-field imaging data from the DESI Legacy Survey combined with a color-magnitude matched-filter technique, we performed a systematic analysis of extratidal features in 28 Galactic globular clusters of likely extragalactic origin, representing the largest homogeneous sample examined in this context to date. The clusters display diverse morphologies: 12 exhibit tidal tails, nine show diffuse envelopes, and seven reveal no clear extratidal features. Notably, we report the first detection of an extended tidal structure around the Sagittarius-associated cluster Terzan 7. To explore the underlying drivers, we compared intrinsic properties, orbital dynamics, and possible accretion associations across morphological groups. From the parameter distributions, complemented by Kolmogorov-Smirnov tests, we find that total mass, escape velocity, concentration, tidal filling factor, pericentric radius, eccentricity, and radial angle in action-angle coordinates are all likely group-sensitive parameters. These results suggest that both internal cluster properties and orbital configurations play important roles in shaping extratidal morphologies. In addition, the cluster’s accretion history shows no clear correlation with the presence of tidal features, indicating that it is not a direct driver of outer structure formation. Overall, the diversity of tidal structures is unlikely to be governed by a single factor. Instead, it reflects the interplay between internal dynamical evolution and the external Galactic environment. This study provides the most comprehensive constraints so far on the physical processes driving extratidal structures in accreted globular clusters.
@@ 1 Introduction
Globular clusters (GCs) are among the oldest stellar systems in the Universe, and they serve as valuable tracers of the formation and assembly history of their host galaxies. In the hierarchical framework of Lambda cold dark matter cosmology model (Planck Collaboration et al. 2016), many Galactic GCs are thought to have been accreted from satellite galaxies during past merger events (Searle and Zinn 1978; Bullock and Johnston 2005; Kruijssen et al. 2020). This idea is supported by the discovery of phase-space substructures and by the association of certain clusters with known merger remnants, such as the Sagittarius dwarf galaxy or Gaia-Sausage-Enceladus (see, e.g., Law and Majewski 2010; Sohn et al. 2018; Helmi et al. 2018; Massari et al. 2019, and references therein). These GCs of extragalactic origin are believed to have complex merger histories and dynamical perturbations, which may render them more susceptible to developing prominent extratidal features under the influence of tidal forces (Searle and Zinn 1978; Bernard et al. 2016; Grillmair 2017; Renaud et al. 2017; Carlberg 2018).
Tidal structures are generally believed to form when stars acquire sufficient energy to escape the gravitational potential of their host GCs. In the early stages of cluster evolution, energetic feedback from stellar evolution, gas expulsion, and dynamical disequilibrium can cause young clusters to rapidly lose their gravitational binding and dissolve (Geyer and Burkert 2001; Kroupa 2001; Goodwin and Bastian 2006; Baumgardt and Kroupa 2007). Over longer timescales, stellar escape can be driven by a combination of internal and external mechanisms. Internally, two-body relaxation redistributes energy among stars, enabling some to reach escape velocity (Küpper et al. 2008a, b). Externally, tidal stripping induced by the Galactic potential as well as dynamical perturbations such as disk and bulge shocking during the cluster’s orbital passages through the Galactic plane or central regions can further enhance stellar mass loss and facilitate the development of extratidal features (Vesperini and Heggie 1997; Baumgardt and Makino 2003; Montuori et al. 2007; Hozumi and Burkert 2015).
Due to the complex formation pathways and the diverse Galactic environments in which they reside, GCs exhibit a wide variety of tidal structures, such as symmetric or multi-armed tidal tails and diffuse outer envelopes. However, the physical mechanisms driving the formation of these different tidal structures are still under active investigation. Early numerical N-body simulations by Montuori et al. (2007) suggest that repeated pericentric passages or strong interactions with the dense regions of the Galaxy can trigger the stripping of outer stars, leading to the formation of tidal tails. In contrast, the work by Küpper et al. (2010, 2012) highlights the role of epicyclic motions of escaping stars in shaping the formation and morphology of tidal features. Other studies have pointed out that orbital parameters, such as inclination and eccentricity, play an important role in regulating the mass loss of clusters (Piatti 2019; Piatti and Carballo-Bello 2020), which in turn affects their susceptibility to developing tidal structures. Meanwhile, the potential connection between cluster origin and structural properties has garnered increasing attention. For example, whether the presence of tidal structures is connected to specific accretion events, such as the Sagittarius dwarf galaxy or the Gaia-Sausage-Enceladus merger, remains to be systematically investigated.
Benefiting from the advent of wide-field imaging surveys such as the Sloan Digital Sky Survey (SDSS; Abazajian et al. 2009), Pan-STARRS (Chambers et al. 2016), the Dark Energy Survey (DES; Abbott et al. 2018), and the DESI Legacy Imaging Surveys (Dey et al. 2019), together with the development of various techniques optimized for identifying tidal substructures, most notably the application of matched filter method, previous studies have uncovered an increasingly rich diversity of tidal features, including long tidal tails (e.g., Palomar 5 and NGC 5466; Grillmair and Dionatos 2006; Grillmair and Johnson 2006), multi-arm structures (e.g., NGC 4147; Jordi and Grebel 2010; Zhang et al. 2024), and a variety of other complex substructures reported in the literature (Lauchner et al. 2006; Olszewski et al. 2009; Chun et al. 2010; Jordi and Grebel 2010; Carballo-Bello et al. 2012, 2014; Kuzma et al. 2016; Carballo-Bello et al. 2017; Navarrete et al. 2017; Carballo-Bello et al. 2018; Kuzma et al. 2018; Muñoz et al. 2018b; Musella et al. 2018; Shipp et al. 2018; Carballo-Bello 2019; de Boer et al. 2019; Grillmair 2019; Palau and Miralda-Escudé 2019; Piatti and Fernández-Trincado 2020; Shipp et al. 2020; Sollima 2020; Thomas et al. 2020; Ibata et al. 2021; Piatti et al. 2021; Nie et al. 2022; Piatti 2022; Zhang et al. 2022; Awad et al. 2025; Chiti et al. 2025; Wan et al. 2025; Wang et al. 2025; Yang et al. 2025). Precisely because such a large sample of GCs with diverse tidal morphologies has now become available, we are afforded the opportunity to investigate, from an observational perspective, the key physical mechanisms that drive the diversity of tidal structures in these systems.
Recent studies have utilized observational data to explore the key factors that drive the formation of different types of tidal structures. Piatti and Carballo-Bello (2020) conducted a multi-parameter analysis of 53 Galactic GCs compiled from previous literature, including both in situ and accreted systems. However, their classification of tidal structures relied on results from multiple studies that used heterogeneous datasets obtained with different telescopes. More recently, Kuzma et al. (2025) constructed a smaller sample of 30 GCs from the Pristine-Gaia-Synthetic catalog, likewise covering both in situ and accreted clusters, and in a brief section of their work, they also performed an exploratory parameter correlation analysis. These studies have provided valuable attempts to explore possible connections between cluster properties and tidal structures, although no significant correlations were found. Nevertheless, studies based on homogeneous data that focus exclusively on a specific class of globular cluster systems, such as those accreted from outside, are still lacking. Such clusters, owing to their external origin, may have undergone distinct pathways of dynamical evolution and tidal disruption (Massari et al. 2019; Forbes 2020).
The objective of this paper is to investigate the key factors that give rise to the tidal structures of the accreted globular cluster population by employing a homogeneous dataset and a unified methodology. We used the deepest dataset among the currently available public data and the largest accreted cluster sample among those widely recognized (Massari et al. 2019; Forbes 2020). Combined with a consistent matched-filter technique, this study minimizes the methodological biases and systematic uncertainties.
This paper is organized as follows. In Sect. 2 we present the selection criteria for the globular cluster sample and introduces the photometric data employed. In Sect. 3 we outline the matched-filter technique adopted to search for tidal tails around GCs. In Sect. 4 we present the main results, including the identification and classification of tidal structures, their correlation with physical parameters, orbital properties, and the Galactic environment, as well as a discussion of the limitations in detecting such structures. In Sect. 5 we summarize the main findings of this study.
2 Observational material
We began by constructing an initial list of candidate GCs in the Galaxy with probable extragalactic origins based on consistent classifications in both Massari et al. (2019, 2025 update) and Forbes (2020). In the original 2019 version, Massari et al. (2019) classified clusters primarily according to their positions in integrals-of-motion (IOM) space derived from Gaia DR2 astrometry, with the age-metallicity relation serving as a secondary criterion. In the 2025 update, this classification was refined using Gaia eDR3 astrometry together with two independent sets of distance estimates from Baumgardt and Vasiliev (2021) and Harris (1996, 2010 edition). We retained only those clusters that were consistently classified as accreted under both distance assumptions. In contrast, Forbes (2020) adopted the IOM framework of Massari et al. (2019) but also incorporated the updated compilation of cluster ages and metallicities from Kruijssen et al. (2019), increasing the number of clusters with complete chemical information. Where available, they included $\alpha$-element abundances from Horta et al. (2020), enabling more reliable assignments of clusters to disrupted dwarf galaxies and, in some cases, revising the original classifications. By cross-matching the results of these two studies and excluding clusters with unclear or ambiguous origins, we obtained an initial sample of 44 GCs of extragalactic origin.
@@ 2 Observational material
We selected stellar sources classified as point-source in the DESI Legacy Surveys catalogs, requiring $g\leq 24$ and $r\leq 23$, with photometric uncertainties of less than 0.2 mag in both bands, in order to ensure a more complete and reliable star sample. For clusters lacking $r$-band data (e.g., Ter 7, Ter 8, and Pal 12), we substituted $i$-band photometry and applied the same selection criteria to maintain consistency. In the DESI Legacy Surveys catalogs, some photometric measurements (such as those in the $i$-band) are drawn from supplementary datasets incorporated into the survey products (e.g., the DES, the DELVE Survey, and the DeROSITA Survey). Subsequently, all magnitudes were corrected for Galactic extinction using the dust maps of Schlegel et al. (1998), recalibrated by Schlafly and Finkbeiner (2011), with extinction coefficients of 3.214, 2.165, 1.592, and 1.211 for the $g$, $r$, $i$, and $z$ bands, respectively, following the calibrations of Dey et al. (2019).
@@
Finally, we examined whether each cluster in the initial sample lies within the footprint of the DESI Legacy Imaging Surveys and possesses sufficient photometric coverage. Among them, 11 clusters (NGC 2298, NGC 3201, NGC 4833, NGC 5286, NGC 6235, NGC 6333, IC 1257, FSR 1758, ESO-SC06, NGC 6715, and NGC 6779) are located in regions not covered by DESI Legacy Surveys. In addition, NGC 2808 and NGC 6101 are covered only in the $g$ and $i$ bands, respectively; NGC 5139 suffers from severely incomplete coverage, with many areas lacking photometric detections; IC 4499 is heavily contaminated by field stars and has poor photometric quality; and NGC 2419, at a distance of about 88 kpc (Baumgardt and Vasiliev 2021), is too remote for our photometric selection criteria to be reliably applied. After excluding these systems, we obtained a final working sample of 28 clusters (see Table 1). These include clusters with well-established progenitors (such as Gaia-Sausage-Enceladus (GSE), Sagittarius (Sag), Sequoia (Seq), and the Helmi streams (H99)) as well as clusters of clear accreted origin but uncertain association with a specific progenitor. Although some previous compilation studies have reported a comparable number of accreted GCs, the sample constructed in this work is entirely based on uniformly deep imaging data from the DESI Legacy Surveys, making it the deepest and largest homogeneous dataset to date for studying the tidal structures of GCs of extragalactic origin within a consistent data framework.
@@
Table 1: Internal parameters and classification of our globular cluster sample. [^1] Columns list: cluster name, King and dynamical tidal radii, present-day cluster mass (logarithmic scale), central escape velocity, central concentration, ratio of half-mass radius to dynamical tidal radius, mass-loss fraction, metallicity, relaxation-age ratio, classification type, and possible progenitor system. Among them, $r_{\rm t}^{\rm King}$, $c$ and [Fe/H] are taken from the Harris catalog (Harris 1996, 2010 edition); the classification type and progenitor system are derived in this work; all other parameters are taken from, or computed using, the globular cluster database used in this study.
@@ 3 Matched filter method
To detect potential extratidal structures around each globular cluster, we employed the matched filtering technique developed by Rockosi et al. (2002), which enhances the contrast between cluster members and field stars in CMD space and is effective for detecting low-surface-brightness structures. This technique has been successfully applied in several subsequent studies of tidal features around star clusters (e.g., Chun et al. 2010; Jordi and Grebel 2010; Navarrete et al. 2017; Carballo-Bello et al. 2018; Shipp et al. 2018; Sollima 2020; Nie et al. 2022; Yang et al. 2023; Zhang et al. 2024), demonstrating its effectiveness in tracing surrounding stellar substructures.
According to Rockosi et al. (2002), at any position on the sky (within a solid angle $d\Omega$), the expected number of stars in any sky bin $(i,j)$ in color-magnitude can be expressed as:
@@
Here, $f_{\mathrm{cl}}$ represents the normalized color-magnitude distribution of cluster members and serves as the cluster template in the matched-filter method, while $n_{\mathrm{bg}}$ denotes the background number density distribution in the same bin, which is used as the corresponding background template. The parameter $\alpha$ serves as a scaling factor that quantifies the number (or surface density) of stars following the cluster’s CMD distribution within the given sky region. Subsequently, the calculation of this parameter can be derived using the minimum-variance estimator (see Equation 2 in Rockosi et al. 2002).
@@
Based on this model, we first constructed the cluster CMD template ($f_{\mathrm{cl}}$) used in the matched-filtering analysis. To do so, we generated a Hess diagram for the cluster region using bin sizes of $\Delta(g-r)=0.05$ and $\Delta r=0.1$ mag. In most cases, the CMD was built in the (gr,r)(g-r,r) plane. However, for a few clusters lacking $r$-band photometry, such as Ter 7, Ter 8, and Pal 12, the (gi,i)(g-i,i) combination was adopted instead. In order to reduce the effects of crowding and incompleteness in the cluster core, stars within the inner 1′-3′ were excluded from the template construction. For clusters with relatively small tidal radius and more diffuse stellar distributions (e.g., Whiting 1 and Pal 13), no such exclusion was applied. Consequently, the cluster template was constructed from stars located between this exclusion zone (when applied) and the cluster’s tidal radius (from Harris (1996), 2010 edition). In addition, to reduce field-star contamination and obtain a clean cluster CMD template $f_{\mathrm{cl}}$, we selected stars with 0.5<gr<1.0-0.5<g-r<1.0 and $14<r<23$ mag, and subtracted a local background derived from a comparing field in an annular region centered on the cluster, typically located between about $1^{\circ}$ and $1.5^{\circ}$ from the cluster center. The stellar density in this annular region was normalized to the area of the cluster region and subtracted from the cluster Hess diagram, yielding a background-subtracted CMD template (see the upper panels of each Hess diagram in Fig. 1). In contrast, the global background CMD template ($n_{\mathrm{bg}}$) required for the matched-filtering analysis was constructed from four rectangular control regions ($1^{\circ}\times 1^{\circ}$) located at different position angles far away from the cluster, each more than $2^{\circ}$ from the cluster center. The Hess diagrams of these regions were then averaged to produce a uniform global background model (see the lower panels of each Hess diagram in Fig. 1).
@@
Finally, the matched-filter procedure yields the quantity $\alpha$, which represents the number of stars that pass the matched filter within a given spatial bin. Using these $\alpha$ values, we then constructed a spatial density map to reveal possible tidal structures. However, if the spatial bins are too large, thin tidal features may be smeared out, whereas overly small bins introduce noise and may generate spurious features. Therefore, after testing a range of spatial binning scales and visually inspecting the results, we ultimately adopted a uniform grid with an angular resolution of $0.02^{\circ}\times 0.02^{\circ}$. To evaluate the statistical significance of the detected structures, we computed a signal-to-noise (S/N) map. For each spatial bin, we subtracted the mean background level (estimated from the outer regions excluding the cluster) from its $\alpha$ value and then divided the result by the standard deviation of the background. This final S/N map reveals potential extratidal features with enhanced contrast (see Fig. 2).
@@
Figure 1: Color-magnitude diagrams for part of the cluster sample. Each panel displays the template Hess diagram (top) and the global background Hess diagram (bottom). CMDs for the remaining clusters are presented in Fig. 7.